# An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors

### v3

@@ 4.3.1. Model Stability
To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
@@
Table 7. Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
@@
4.3.2. Explanatory Power of Each Factor
@@
4.3.3. Relationships between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 7). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files that have fewer tokens, contain functional logic, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
5. Discussion
@@
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (Steenhoek et al., 2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy. However, in real-world security reviews, although with the experience and domain knowledge, The reviewers lack prior knowledge of all potential types of security defects that may be present in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects.
@@
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better-supporting LLMs in code review tasks.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
@@
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because the outputs generated by LLMs with higher randomness are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower randomness. Both detection capability and the stability of detection results are crucial for the practical application of LLMs in security code review. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the randomness inherent in LLMs represents a valuable direction for future improvements.
6. Threats to Validity
@@
Internal Validity. One threat could be data leakage. The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs. Since the training data cutoff dates for GPT-3.5 and GPT-4 are not disclosed and the dataset we used (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to the findings by Cao et al. (Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To specifically analyze its impact, we designed prompt $\textbf{P}_{cot}$, in which we instructed LLMs to generate code context according to the commit message before detecting security defects. Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness. To mitigate this threat, we conducted three repetitive experiments for RQ1. For RQ2, we calculated the average frequency of each problem across these three experiments. In RQ3, we combined the responses from the three experiments to fit the model.
Conclusion Validity. In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale). In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies. We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
External Validity. With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure. These models are also not made publically available via API, thereby hindering the reproducibility of this study. Therefore, we explicitly indicate the versions of these two models used in the experiments and provide all the responses generated by these models (Yu et al., 2024). In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements. Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
7. Conclusions and Future Work
@@
In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM. We performed a regression analysis to explore factors influencing its performance.
@@
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v4

@@ 3.3.1 Model Stability
To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the non-determinism of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
@@
Table 8: Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
@@
3.3.2 Explanatory Power of Each Factor
@@
3.3.3 Relationships between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 8). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files with fewer tokens, that contain functional logic, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that annotations may disrupt the LLM and are less likely to detect the target security defect.
4 Discussion
@@
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy. However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al. (2023), providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects. Due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible. The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version. We tested various approaches, including one-shot and few-shot prompting strategies, as well as providing the complete patchset within the prompt. These strategies required splitting the code into sections for input, which we found significantly degraded the performance of LLMs compared to the basic prompt $\textbf{P}_{b}$. In order to provide as much effective information as possible within the limited context window, we designed prompts that combined multiple elements, including project metadata, CWEs, code diffs, and the commit messages. We also standardized the output format to help the model better understand the task and introduced CoT reasoning to guide the model in inferring missing context. Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments. The results demonstrate that including a list of CWEs significantly enhances LLM performance, making it our study’s most effective prompting strategy. As the context window of LLMs continues to expand and prompt engineering techniques evolve, we plan to explore additional strategies in the future, such as incorporating RAG for external information and employing automated prompt optimization to dynamically adjust prompts during reasoning, thereby further improving the capability of LLMs in security defect detection.
@@
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better supporting LLMs in code review tasks.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. Therefore, ensuring the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
@@
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because the outputs generated by LLMs with higher non-determinism are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower non-determinism. Both detection capability and the stability of detection results are crucial for applying LLMs in security code review. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the non-determinism inherent in LLMs represents a valuable direction for future improvements.
5 Threats to Validity
@@
Internal Validity refers to whether an observed association between two variables can be attributed to a causal link between them. One of the internal validity threats could be data leakage. The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies. Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to Cao et al. (2024), contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, non-suggestive, and unambiguous as possible. We iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To specifically analyze its impact, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to generate code context according to the commit message before detecting security defects.
External Validity concerns whether a causal link generalizes across contexts. In this study, the experiments were conducted on 6 LLMs based on 4 open-source projects across programming languages including Python and C/C++. We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs. We implemented the following strategies to ensure the generalizability of the experimental results. For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular LLMs. For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities. These projects differ in architectures, programming languages, and community structures, providing a comprehensive and valuable dataset for empirical analysis.
Construct Validity assesses whether a specific set of metrics corresponds to what they are intended to measure. To measure the performance of LLMs, we classified the natural language responses of LLMs into four predefined categories, with human reviews of security defects in code as the ground truth. There is a potential natural language understanding bias which affects classification accuracy. To reduce the bias, we established granular definitions for each category in three dimensions: defect type, location, and description, supplemented by examples for each category to ensure that the metrics are accurate, appropriate, and consistent throughout the evaluation process. In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually. In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (Glaser, 1965). In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion. In the data analysis of RQ2, the first and seventh authors, conducted two rounds of pilot data analysis to determine the definition of each quality problem and achieved a high level of consistency in the results. All discrepancies in the manual work of our study were reviewed and resolved through negotiation and discussion among at least three authors, thus ensuring the reliability of the findings.
Conclusion Validity refers to whether the conclusions drawn about relationships in the data are statistically sound and reasonable. Due to the non-determinism of LLMs, the same input may yield different outputs across different runs. Therefore, a single round of experiments is insufficient to accurately reflect the performance of LLMs. To enhance the reliability and representativeness of our evaluation, we repeated the RQ1 experiment three times. For RQ2, we calculated the average frequency of each problem across these three experiments. For RQ3, we combined the responses from the three experiments to fit the model. We also assessed their non-determinism as part of the overall performance. Specifically, in RQ1.4, we calculated the consistency of LLMs’ responses across three runs to compare the non-determinism of different LLMs under different prompts.
6 Related Work
6.1 Security Defect Detection in Code Review
Several studies have analyzed security defect detection during code reviews. Paul et al. (2021) built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects. The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects. It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact. Alfadel et al. (2023) examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews. Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
6.2 Code Characteristics of Security Defects
Security defects mainly stem from improper system design or low-quality code, caused by a lack of security expertise or violations of secure coding practices during implementation. Many studies have focused on analyzing code for vulnerabilities and security defects in hopes of helping developers understand how security problems originate in a system, so that defects are detected early.
Alves et al. (2016) constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others. This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners. Meneely et al. (2013) conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects. The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects. Bosu et al. (2014) identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with Meneely et al. (2013) that longer lines of code and less experienced committers are more likely to introduce security defects. Additionally, they indicated that modified code files are more likely to contain security defects than new files.
6.3 Automated Tools of Security Defect Detection
Various tools have been developed to automatically detect security defects. Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or program repair.
Charoenwet et al. (2024) conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs. By combining several tools, warnings could be produced for 78% of the VCCs. However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review. Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. (2024b) utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts. They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. (2023) compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools. Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. (2023) provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers. Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin and Ni (2024) evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed LLMs. Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks. These studies have all been limited to quantitatively evaluating LLM performance. In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
Employing specific prompting frameworks to enhance the performance of LLMs in identifying security defects has also been explored. Yang et al. (2024) introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection. Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. (2024a) proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution). The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours. These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements. Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs. Inspired by these related works, we aim to leverage LLMs’ potential in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
7 Conclusions and Future Work
@@
In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM. We performed a regression analysis to explore factors influencing its performance.
@@
Data availability
We have shared the link to our dataset in the reference (Yu et al., 2025).
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.
[…]